Simplify the answer, round to the nearest hundreth, and type an integer or decimal.
QUESTION:
During the first part of a trip, a canoeist travels over 77 miles at a certain speed. The canoeist travels 11 miles on the second part of the trip at a speed of 5mph slower . The total time for the trip is 3 hours. What was the speed on each part of the trip?

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



WRONG

If the canoeist used the ENTIRE 3 hours to complete the first portion of the trip, s/he would have to make the canoe move at a rate in excess of 25 miles per hour. Not happening. The world record for a 1000 meter sprint is less than 11 miles per hour. Check your numbers and repost.


John


Answer by stanbon(75887)

During the first part of a trip, a canoeist travels over 77 miles at a certain speed. The canoeist travels 11 miles on the second part of the trip at a speed of 5mph slower . The total time for the trip is 3 hours. What was the speed on each part of the trip?
------------------------------
1st Part DATA:
distance = 77 miles ; rate = x mph ; time = 77/x hrs
----------------------
2nd Part DATA:
distance = 11 miles ; x-5 mph ; time = 11/(x-5) hrs
-----
Equation:
time + time = 3 hrs
77/x + 11/(x-5) = 3
77(x-5) + 11x = 3x(x-5)
77x - 385 + 11x = 3x^2 - 15x
3x^2 - 103x + 385 = 0
x = 4.2686 or x = 30.06
---
Since x-5 would be negative if x = 4.2686,
the only solution is x = 30.06 mph
---
1st part rate = 30.06 mph
2nd part rate = 25.06 mph
==========================
Cheers,
Stan H.
==========================